OpenStudy (anonymous):

A coin is unbalanced such that it comes up tails 55% of the time. a) What is the probability that you don't flip a head until the 4th flip? b) What is the probability that you don't flip a tail until the 4th flip? c) How many tosses would it take, on average, to flip a head? d) How many tosses would it take, on average, to flip a tail?

OpenStudy (anonymous):

@robtobey what do you think

OpenStudy (anonymous):

a. \[(55/100)^3 (45/100) \]

OpenStudy (anonymous):

for a i get .07486875

OpenStudy (anonymous):

0.07487 is what I got after rounding.

OpenStudy (anonymous):

okay great! wouldn't B be the same answer?

OpenStudy (michele_laino):

I think: \[\large {\left( {\frac{{45}}{{100}}} \right)^3} \times \frac{{55}}{{100}}\]

OpenStudy (anonymous):

so for b) .05011875?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

what about c and d?

OpenStudy (anonymous):

wait I'm confused how do we figure that then

OpenStudy (michele_laino):

I give you more explanation

OpenStudy (michele_laino):

we have a coin and the probability to get a tail is 55/100, whereas the probability to get a head is 45/100 so I call p=45/100 and q=1-p=55/100

OpenStudy (michele_laino):

Now the probability to get a tail after N tosses, is: \[\large \left( {\begin{array}{*{20}{c}} N \\ 1 \end{array}} \right){\left( {\frac{{45}}{{100}}} \right)^{N - 1}}\frac{{55}}{{100}}\]

OpenStudy (michele_laino):

whereas the probability to get a head after N tosses, is: \[\large \left( {\begin{array}{*{20}{c}} N \\ 1 \end{array}} \right){\left( {\frac{{55}}{{100}}} \right)^{N - 1}}\frac{{45}}{{100}}\]

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

so, for answer c and d, we have to compute such N in order to get the requested probability

OpenStudy (anonymous):

how do we get N

OpenStudy (michele_laino):

we can get N using the formulas above.

OpenStudy (michele_laino):

please wait, I'm confused, because, if I apply the binomial distribution, then the answers to questions a. and b. would be different by a factor which is a binomial coefficient

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

I ask to another tutor, please wait... @dan815 can you help here please?

OpenStudy (michele_laino):

@TuringTest can you help here please?

OpenStudy (michele_laino):

@amistre64 can you help here please?

OpenStudy (turingtest):

You don't need the binomial coefficient

OpenStudy (turingtest):

you would need the binomial coefficient if the question was "what are the odds of getting one heads/tails within the first 4 tosses"

OpenStudy (michele_laino):

ok! what are the answers to questions c. and d. @TuringTest

OpenStudy (amistre64):

hmm, if we wanted to know the number .. yeah

OpenStudy (turingtest):

There is a nifty proof to the effect that a geometric rv with probability p has expected value of 1/p. I would be hard pressed to derive it off the top of my head though :P

OpenStudy (amistre64):

P(1h in 4 tosses) = (4 1) ... ut not getting it til the 4th toss is just: P(t)P(t)P(t)P(h)

OpenStudy (amistre64):

How many tosses would it take, on average, to flip a head? h = 45 out of 100 45/45 = 100/45 seems reasonable to me

OpenStudy (anonymous):

so 45/45 is the answer?

OpenStudy (amistre64):

of course 45/45 is not the answer ...

OpenStudy (anonymous):

im so confused

OpenStudy (michele_laino):

so you are referring to the mean value and to this equation: 1 = N*p 1= N*(45/100)? @amistre64

OpenStudy (amistre64):

45 tosses out of 100 tosses on average what is 1 toss out of n tosses?

OpenStudy (anonymous):

2.222222

OpenStudy (michele_laino):

I have thought to that formula, but I was not sure

OpenStudy (amistre64):

i have no reference, just a little common sense is all say its fair 50h out of 100 tosses 50/50 = 100/50 1h in 2 tosses

OpenStudy (turingtest):

Here is proof that amistre is right http://mathaa.epfl.ch/cours/PMMI2001/interactive/geomexpect_en0.htm

OpenStudy (amistre64):

45 h out of 100 tosses, we woud expect to get 45/45 heads out of 100/45 tosses

OpenStudy (michele_laino):

ok! thank you! @amistre64 @TuringTest

OpenStudy (amistre64):

since 2.22 is not a valid toss ... 3 tosses should suffice

OpenStudy (anonymous):

okay so these are my answers: a) .07486875 b).05011875 c) 2.222 rounded to three d) 1.818 rounded to two

OpenStudy (amistre64):

seems good to me yes

OpenStudy (anonymous):

okay thank you all so much!

OpenStudy (michele_laino):

thank you all!

OpenStudy (amistre64):

yalls welcome :)